OpenStudy (anonymous):

Bert, Burt, and Belle are professional elephant manicurists. Bert can manicure an elephant toe in 5 minutes, Burt can manicure an elephant toe in 8 minutes, and Belle can manicure an elephant toe in 5 minutes. How long, to the nearest minute, would it take the three of them working together to manicure five elephant toes? Please show me your work.

hero (hero):

you posted the same time twice

hero (hero):

Bert and Belle can both manicure an elephant toe in 5 minutes...are you sure that's right?

OpenStudy (anonymous):

yeah! I am sure

hero (hero):

\[\frac{1}{5}+\frac{1}{8}+\frac{1}{5} = \frac{5}{x}\] \[\frac{2}{5}+\frac{1}{8}=\frac{5}{x} \] Multiply both sides by x: \[\frac{2x}{5}+\frac{x}{8} = 5\] Change 5 into a fraction: \[\frac{2x}{5}+\frac{x}{8} = \frac{40}{8} \] Subtract x/8 from both sides: \[\frac{2x}{5} = \frac{40}{8}-\frac{x}{8}\] Simplify right side: \[\frac{2x}{5}=\frac{40-x}{8}\] Cross Multiply: 16x = 5(40-x) 16x = 200 - 5x 16x+5x = 200 21x = 200 x = 9.5 approximately 10

OpenStudy (anonymous):

Thanks